Determine the area of the regular octagon circumscribing a circle having an area of 126 m2.
A. 127.83 m2
B. 132.90 m2
C. 119.05 m2
D. 113.44 m2

Answer by Boreal(15235)

Draw this
The area of the circle is 126 m^2=pi*r^2
therefore, r=6.333
Each of the 8 parts of the octagon is a triangle with central angle 45 deg
Draw a perpendicular to the side, and this bisects the central angle.
That triangle has angle 22.5, the opposite side is half the length of a side of the octagon, and the perpendicular line is the adjacent side.
cos22.5=adj/6.333; adj side =5.851
sin 22.5=opp/6.333, opp side=2.424
the side of the octagon is 4.848, double the above for sin 22.5
the altitude is 5.851
1/2 that product is the area of the triangle.
There are 8 such triangles, so 4 times the product is the answer or 113.46 m^2.
D.